# Understanding Figurative Meaning through Explainable Visual Entailment

### v2

@@ Abstract
Large Vision-Language Models (VLMs) have demonstrated strong capabilities in tasks requiring a fine-grained understanding of literal meaning in images and text, such as visual question-answering or visual entailment. However, there has been little exploration of these models’ capabilities when presented with images and captions containing figurative meaning, such as metaphors or humor. To close this gap, we propose a new task framing the figurative meaning understanding problem as an explainable visual entailment task, where the model has to predict whether the image (premise) entails a caption (hypothesis) and justify the predicted label with a textual explanation. The figurative phenomena can be present either in the image, the caption, or both. Utilizing a human-AI collaboration approach, we build the accompanying expert-verified dataset V-FLUTE, containing 6,027 image, caption, label, explanation instances spanning five diverse figurative phenomena: metaphors, similes, idioms, sarcasm, and humor. Through automatic evaluation, we find that VLMs struggle to generalize from literal to figurative meaning, particularly when it is present in images. Further, we identify common types of errors in VLM reasoning via human evaluation.111Code and data: github.com/asaakyan/V-FLUTE
@@ 1 Introduction
However, figurative meaning is also prevalent in visual phenomena, such as visual metaphors Akula et al. (2023); Chakrabarty et al. (2023), multimodal sarcasm Desai et al. (2022), and humor Hessel et al. (2023); Hwang and Shwartz (2023). Yet so far most of the work on vision and language models (VLMs) has focused on understanding literal meaning in images and captions (e.g., ScienceQA Lu et al. (2022), MMMU Yue et al. (2024)) including work on explainable visual entailment Kayser et al. (2021). Building on the idea of FLUTE Chakrabarty et al. (2022) for text, we present a new dataset for understanding figurative meaning as explainable visual entailment, we call V-FLUTE. Our dataset contains 6,027 image, caption, label, explanation instances spanning diverse figurative phenomena. Each instance contains an image (premise) and a caption (hypothesis) that is either entailed or contradicted by the image. Deciding the entailment relation requires the vision-language model to understand the implicit meaning in both the visual and textual modalities. Our dataset contains figurative phenomena present in the image, in the caption, or in both. In addition, to mitigate the dependence on spurious correlations, to more rigorously investigate reasoning capabilities, and to promote explainability, our task requires the model to generate a plausible explanation for the output label. See Figure 1 for two examples from our dataset.
@@ 3 V-FLUTE Task and Dataset
To build V-FLUTE, we start with existing multimodal figurative datasets which cover phenomena such as metaphors, similes, idioms, sarcasm or humor. We utilize human-AI collaboration frameworks with expert annotators Chakrabarty et al. (2022); Wiegreffe et al. (2022); Liu et al. (2022) to augment them with expert-verified textual explanations and entailing/contradicting captions. Each instance then includes an image and a caption, and the figurative phenomenon can be either in the image, the caption or in both. An overview of V-FLUTE dataset and our contributions w.r.t to the source datasets can be found in Table 1. See examples corresponding to each source dataset in Table 2 as they appear in V-FLUTE. Below, we describe the construction of V-FLUTE by each phenomenon.
@@ 3.1.1 IRFL as Data Source
Expert Verification. We recruit three expert annotators with significant experience in figurative language and visual metaphor understanding on Upwork and ask them to verify the explanation is correct, complete, and concise and if not, edit it (see details in Appendix B). We also ask the annotators to discard rare noisy instances where the caption, image, and label do not fit (due to automatic image selection). Due to relative simplicity of generating the explanation given a literal image, the experts only needed to edit % of the explanations. They also removed the data, resulting in 1149 image, caption, label, explanation instances for metaphors and similes and 370 for idioms.
@@ 3.1.2 HAIVMet as Data Source
Chakrabarty et al. (2023) use a human-AI collaboration framework to generate visual metaphors from linguistic metaphors (HAIVMet dataset) and propose a visual entailment task as an extrinsic evaluation of dataset quality. The HAIVMet data consists of 1,193 images of visual metaphors spanning over 958 distinct linguistic metaphors. Each image is associated with a caption that can be contradicting or entailing the image. In addition, each image is associate with a visual elaboration that presents a textual description of the image (See Appendix A, Figure 7). This visual elaboration was used in the original paper to generate the visual metaphors (images).
Generating Textual Explanations. We augment the dataset with candidate textual explanations. We prompt ChatGPT (gpt-3.5-0914) to generate an explanation for every tuple visual elaboration, caption, label (See Appendix A, Figure 7; and prompt in Appendix F.1.1).
Expert Verification. Each caption is paired with up to images. However, since these images were automatically generated with DALLE-2 using the visual elaborations, not all are completely faithful. Moreover, some captions and labels were inconsistent. Finally, automatically generated LLM candidate explanations are not always correct and require refining. To tackle these issues, we employ an expert verification process recruiting the same three expert annotators as from the IRFL section above (see details in Appendix B). We ask the annotators to select the visual metaphor most faithful to the linguistic metaphor and the visual elaboration (see Image Selection in Appendix A, Figure 7) or if none were. In addition, we ask them to verify and edit the explanation if necessary to ensure correctness, completeness, and conciseness. On average, experts edited of the explanations and of captions, and rejected of visual metaphors, resulting in 857 image, caption, label, explanation instances.
@@ 3.2.1 MuSE as Data Source
Generating Textual Explanations. While the dataset already contains crowdworker-written explanations, upon inspection, they were often deemed poor quality, lacking enough details, and formulaic (e.g., see the crowdworker explanation in Figure 3). To improve their quality, we use the dataset’s existing crowdworker explanations and prompt GPT-4 to rewrite and generate candidate textual explanations given the caption and the label (see the re-written explanation in Figure 3). See the prompt in Appendix F.3.
@@ 3.3.1 MemeCap as Data Source
Generating Textual Explanations. We prompted GPT-4 with the ground truth label in the prompt to explain the relationship between the image and the caption. See prompts in Appendix F.4.
@@ 3.4 Dataset Statistics
We split our data into 4,578 training, 726 validation, and 723 testing instances. Detailed counts per phenomenon and dataset, as well as other statistics, are in Appendix C.
@@ 4 Experiments
We empirically study how several baseline models perform on the task of explainable visual entailment. We investigate both off-the-shelf and fine-tuned model performance. We provide human baseline performance in Appendix K. Hyperparameters are provided in Appendix E.
@@ 4.1 Models
We select a variety of models for our study (see taxonomy in Appendix, Figure 10). For off-the-shelf models, we explore both open and API-based models. For open models, we select the (current) state-of-the-art LLaVA-1.6 models Liu et al. (2024). LLaVA is one of the simplest, yet one of the most high-performing VLM architectures currently available. It utilizes a pretrained large language model (e.g., Mistral-7B Jiang et al. (2023)) and a vision-language cross-modal connector (e.g., an MLP layer) to align the vision encoder (e.g., CLIP Radford et al. (2021)) outputs to the language models. We select LLaVA-1.6 models in their 7B and 34B configurations (LLaVA-v1.6-7B and LLaVA-v1.6-34B respectively) and refer to them as LLaVA-ZS-7B and LLaVA-ZS-34B. Both models have been instruction-tuned on less than 1M visual instruction tuning samples to act as general language and vision assistants. We also utilize Compositional Chain-of-Thought Prompting proposed by Mitra et al. (2023) denoted by LLaVA-ZS-7B-SG and LLaVA-ZS-34B-SG (see description and results discussion in Appendix H).
@@ 4.2 Automatic Metrics
Since our goal is to ensure models provide an answer for the right reasons, ideally, we would only count predictions as correct when the explanation is also correct. Based on prior work Chakrabarty et al. (2022), we use both the standard F1 score and an adjusted score that accounts for explanation quality: F1@ExplanationScore. The ExplanationScore computes the average of BERTScore Zhang* et al. (2020) and BLEURT Sellam et al. (2020) between model-generated and reference (V-FLUTE) explanations. We report F1@0 (simply F1 score), F1@53333Thresholds selected based on human evaluation of explanation quality in Appendix L. (all predictions with ExplanationScore 53 are considered incorrect) and F1@60.
@@ 4.3 Automatic Evaluation Results
We include results per phenomenon in Appendix J, discussion on CoT prompting in Appendix H and additional models in Appendix I. Table 3 shows the results, informing the following insights:
@@ 5.2 What Errors Do Models Make?
6 Conclusions
We introduce a novel dataset for understanding figurative meaning in multimodal input, V-FLUTE, via an explainable visual entailment task. Our dataset consists of 6,027 image, caption, label, explanation instances covering diverse phenomena. We find that VLMs struggle to generalize from literal to figurative meaning, particularly in images, and identify error types in VLM reasoning about multimodal figurative phenomena.
@@ 8 Limitations
Reference-based evaluation has fundamental flaws, such as not considering all possible explanations, which would be impossible to collect. However, current reference-free metrics for free-text rationales may still have flaws such as bias toward length or the evaluator LLM Stureborg et al. (2024); Raina et al. (2024); Huang et al. (2024); Chiang and Lee (2023); Wei et al. (2024). When evaluating textual explanations against these references, as is the case with any reference-based evaluation, there may also be a preference towards models which output text closer in distribution to the GPT-4 model. Because of that, it is important to utilize the data set in order to compare models other than the teacher model and pay more attention to the F1@0 scores, which represent simple classification scores and do not require the outputs to be similar in distribution. In terms of pure F1 score performance, GPT-4 underperforms the fine-tuned model, and performs very closely with Gemini and Claude that were not used to generate the data, with less than 2% difference (see column F1@0, Table 4). Although we showed a relatively high predictive power of automatic explanation scores to predict human judgements (see Appendix L), future work may focus on increasing reliability of reference-based and reference-free textual explanation evaluation methods.
@@
Appendix A Dataset Pipeline Illustrations
We illustrate the pipelines for creating V-FLUTE from each underlying figurative phenomenon in Figures 7, 2, 3, 4.
Appendix B Details on Expert Verification
@@
Appendix C Dataset Statistics
Table 6 shows the number of samples from each source dataset that are included in the randomly selected training, validation, and held-out test splits.
| Type | Dataset | Train | Valid | Test | ||
| Metaphor /Similes | HAIVMET | 649 | 107 | 101 | ||
912 | 117 | 120 | |||
| Idioms | IRFL (idiom) | 170 | 100 | 100 | ||
| Sarcasm | MuSE | 830 | 106 | 106 | ||
| Humor | MemeCap | 1566 | 196 | 196 | ||
| NYCartoons | 451 | 100 | 100 | |||
| Total | 4,578 | 726 | 723 | |||
@@ Appendix C Dataset Statistics
Appendix D API models Hyperparameters
D.1 Claude
@@ D.3 Gemini
Appendix E Fine-tuning Hyperparameters
@@
We utilize LoRA Hu et al. (2022) to fine-tune the models. We utilize the same hyperparameters for all fine-tunes outlined in Appendix E and use early stopping based on a V-FLUTE validation set to prevent overfitting. For evil and e-ViL+V-FLUTE we only fine-tuned for 2 epochs due to size of the e-ViL dataset and took the best checkpoint based on early stopping on V-FLUTE validation set. For eViL we only fine-tuned for 1 epoch to prevent overfitting. For VFLUTE, we trained for 3 epochs, and VFLUTE -Image for 10 epochs (to ensure performance does not increase even with larger number of epochs), for both we took the best checkpoint based on early stopping.
@@
Appendix F Prompts for LLMs
F.1 HAIVMET
F.1.1 One-shot Prompt for generating explanations
@@
F.2 IRFL
F.2.1 Zero-shot Prompt for generating explanations
@@
F.3 MuSE
F.3.1 Few-shot Prompt for generating opposite claims
@@
F.3.2 Zero-shot Prompt for Rephrasing
@@
F.4 MemeCap
F.4.1 Few-shot Prompt for generating entailing claims

### v3

@@ Abstract
Large Vision-Language Models (VLMs) have demonstrated strong capabilities in tasks requiring a fine-grained understanding of literal meaning in images and text, such as visual question-answering or visual entailment. However, there has been little exploration of the capabilities of these models when presented with images and captions containing figurative meaning, such as metaphors or humor. To close this gap, we propose a new task framing the figurative meaning understanding problem as an explainable visual entailment task, where the model has to predict whether the image (premise) entails a caption (hypothesis) and justify the predicted label with a textual explanation. The figurative phenomena can be present in the image, in the caption, or both. Using a human-AI collaboration approach, we build the accompanying expert-verified dataset V-FLUTE, containing 6,027 image, caption, label, explanation instances spanning five diverse figurative phenomena: metaphors, similes, idioms, sarcasm, and humor. Through automatic evaluation, we find that VLMs struggle to generalize from literal to figurative meaning, particularly when it is present in images. Further, we identify common types of errors in VLM reasoning (hallucination and incomplete or unsound reasoning) across classes of models via human evaluation.111Code and data: github.com/asaakyan/V-FLUTE
@@ 1 Introduction
However, figurative meaning is also prevalent in visual phenomena, such as visual metaphors Akula et al. (2023); Chakrabarty et al. (2023), multimodal sarcasm Desai et al. (2022), and humor Hessel et al. (2023); Hwang and Shwartz (2023). Yet so far most of the work on vision and language models (VLMs) has focused on understanding literal meaning in images and captions (e.g., ScienceQA Lu et al. (2022), MMMU Yue et al. (2024)) including work on explainable visual entailment Kayser et al. (2021). Building on the idea of FLUTE Chakrabarty et al. (2022) for text, we present a new dataset for understanding figurative meaning as explainable visual entailment, V-FLUTE. Our dataset contains 6,027 image, caption, label, explanation instances spanning diverse figurative phenomena. Each instance contains an image (premise) and a caption (hypothesis) that is either entailed or contradicted by the image. Deciding the entailment relation requires the vision-language model to understand the implicit meaning in both the visual and textual modalities. Our dataset contains figurative phenomena present in the image, in the caption, or in both. In addition, to mitigate the dependence on spurious correlations, to more rigorously investigate reasoning capabilities, and to promote explainability, our task requires the model to generate a plausible explanation for the output label. See Figure 1 for two examples from our dataset.
@@ 3 V-FLUTE Task and Dataset
To build V-FLUTE, we start with existing multimodal figurative datasets which cover phenomena such as metaphors, similes, idioms, sarcasm or humor. We utilize human-AI collaboration frameworks with expert annotators Chakrabarty et al. (2022); Wiegreffe et al. (2022); Liu et al. (2022) to augment them with expert-verified textual explanations and entailing/contradicting captions. Each instance then includes an image and a caption, and the figurative phenomenon can be either in the image, the caption or in both. An overview of the V-FLUTE dataset and our contributions w.r.t to the source datasets can be found in Table 1. See examples corresponding to each source dataset in Table 2 as they appear in V-FLUTE. Below, we describe the construction of V-FLUTE by each phenomenon.
@@ 3.1.1 IRFL as Data Source
Expert Verification. We recruit three expert annotators with significant experience in figurative language and visual metaphor understanding on Upwork and ask them to verify the explanation is correct, complete, and concise and if not, edit it (see details in Appendix A). We also ask the annotators to discard rare noisy instances where the caption, image, and label do not fit (due to automatic image selection). Due to relative simplicity of generating the explanation given a literal image, the experts only needed to edit % of the explanations. They also removed the data, resulting in 1149 image, caption, label, explanation instances for metaphors and similes and 370 for idioms.
@@ 3.1.2 HAIVMet as Data Source
Chakrabarty et al. (2023) use a human-AI collaboration framework to generate visual metaphors from linguistic metaphors (HAIVMet dataset) and propose a visual entailment task as an extrinsic evaluation of dataset quality. The HAIVMet data consists of 1,193 images of visual metaphors spanning over 958 distinct linguistic metaphors. Each image is associated with a caption that can be contradicting or entailing the image. In addition, each image is associate with a visual elaboration that presents a textual description of the image (See Figure 3). This visual elaboration was used in the original paper to generate the visual metaphors (images).
Generating Textual Explanations. We augment the dataset with candidate textual explanations. We prompt ChatGPT (gpt-3.5-0914) to generate an explanation for every tuple visual elaboration, caption, label (See Figure 3; and prompt in Appendix E.1.1).
Expert Verification. Each caption is paired with up to images. However, since these images were automatically generated with DALLE-2 using the visual elaborations, not all are completely faithful. Moreover, some captions and labels were inconsistent. Finally, automatically generated LLM candidate explanations are not always correct and require refining. To tackle these issues, we employ an expert verification process recruiting the same three expert annotators as from the IRFL section above (see details in Appendix A). We ask the annotators to select the visual metaphor most faithful to the linguistic metaphor and the visual elaboration (see Image Selection in Figure 3) or if none were. In addition, we ask them to verify and edit the explanation if necessary to ensure correctness, completeness, and conciseness. On average, experts edited of the explanations and of captions, and rejected of visual metaphors, resulting in 857 image, caption, label, explanation instances.
@@ 3.2.1 MuSE as Data Source
Generating Textual Explanations. While the dataset already contains crowdworker-written explanations, upon inspection, they were often deemed poor quality, lacking enough details, and formulaic (e.g., see the crowdworker explanation in Figure 4). To improve their quality, we use the dataset’s existing crowdworker explanations and prompt GPT-4 to rewrite and generate candidate textual explanations given the caption and the label (see the re-written explanation in Figure 4). See the prompt in Appendix E.3.
@@ 3.3.1 MemeCap as Data Source
Generating Textual Explanations. We prompted GPT-4 with the ground truth label in the prompt to explain the relationship between the image and the caption. See prompts in Appendix E.4.
@@ 3.4 Dataset Statistics
We split our data into 4,578 training, 726 validation, and 723 testing instances. Table 3 shows the number of samples from each source dataset that are included in the randomly selected training, validation, and held-out test splits. More details in Appendix B.
| Type | Dataset | Train | Valid | Test | ||
| Metaphor /Similes | HAIVMET | 649 | 107 | 101 | ||
912 | 117 | 120 | |||
| Idioms | IRFL (idiom) | 170 | 100 | 100 | ||
| Sarcasm | MuSE | 830 | 106 | 106 | ||
| Humor | MemeCap | 1566 | 196 | 196 | ||
| NYCartoons | 451 | 100 | 100 | |||
| Total | 4,578 | 726 | 723 | |||
@@ 4 Experiments
We empirically study how several baseline models perform on the task of explainable visual entailment. We investigate both off-the-shelf and fine-tuned model performance. We provide human baseline performance in Appendix 5.4. Hyperparameters are provided in Appendix D.
@@ 4.1 Models
We select a variety of models for our study (see taxonomy in Appendix, Figure 10). For off-the-shelf models, we explore both open and API-based models. For open models, we select the (current) state-of-the-art LLaVA-1.6 models Liu et al. (2024). LLaVA is one of the simplest, yet one of the most high-performing VLM architectures currently available. It utilizes a pretrained large language model (e.g., Mistral-7B Jiang et al. (2023)) and a vision-language cross-modal connector (e.g., an MLP layer) to align the vision encoder (e.g., CLIP Radford et al. (2021)) outputs to the language models. We select LLaVA-1.6 models in their 7B and 34B configurations (LLaVA-v1.6-7B and LLaVA-v1.6-34B respectively) and refer to them as LLaVA-ZS-7B and LLaVA-ZS-34B. Both models have been instruction-tuned on less than 1M visual instruction tuning samples to act as general language and vision assistants. We also utilize Compositional Chain-of-Thought Prompting proposed by Mitra et al. (2023) denoted by LLaVA-ZS-7B-SG and LLaVA-ZS-34B-SG (see description and results discussion in Appendix G).
@@ 4.2 Automatic Metrics
Since our goal is to ensure models provide an answer for the right reasons, ideally, we would only count predictions as correct when the explanation is also correct. Based on prior work Chakrabarty et al. (2022), we use both the standard F1 score and an adjusted score that accounts for explanation quality: F1@ExplanationScore. The ExplanationScore computes the average of BERTScore Zhang* et al. (2020) and BLEURT Sellam et al. (2020) between model-generated and reference (V-FLUTE) explanations. We report F1@0 (simply F1 score), F1@53333Thresholds selected based on human evaluation of explanation quality in Section 5.3. (all predictions with ExplanationScore 53 are considered incorrect) and F1@60.
@@ 4.3 Automatic Evaluation Results
We include results per phenomenon in Appendix I, discussion on CoT prompting in Appendix G and additional models in Appendix H. Table 4 shows the results, informing the following insights:
@@ 5.2 What Errors Do Models Make?
5.3 How Well Does the Explanation Score Predict Human Judgment on Adequacy?
We explore whether the proposed explanation score can capture human judgment of explanation adequacy. We collect all instances from Section 5 where both annotators agreed on the adequacy judgement for the explanation. We evaluate if the explanation score described in Section 4.2 can act as a good predictor for the human adequacy judgment. We find that the area under the Precision-Recall curve is 0.79, and the maximum F1 score is 0.77, obtainable at the explanation score threshold of 0.53. Hence, we use this threshold to report the results in Table 4. We also use the threshold of 0.6 since it maximizes F1 such that both precision and recall are above 0.75.
5.4 How Well Do Humans Perform?
| Phenomenon | Dataset |
| Metaphor /Similes | HAIVMET | 78.84 | 81.25 | ||||
94.36 | 77.78 | |||||
| Idioms |
89.26 | 49.74 | ||||
| Sarcasm | MuSE | 68.89 | 85.42 | ||||
| Humor | MemeCap | 100.0 | 78.03 | ||||
| NYCartoons | 71.43 | 47.83 | |||||
| Overall | 89.09 | 77.26 | |||||
6 Conclusion
We introduce a novel dataset for understanding figurative meaning in multimodal input, V-FLUTE, via an explainable visual entailment task. Our dataset consists of 6,027 image, caption, label, explanation instances covering diverse phenomena. We find that VLMs struggle to generalize from literal to figurative meaning, particularly in images. When figurative meaning is present in the image rather than text, models perform worse. VLMs benefit from the visual information during training to understand visual figurative meaning. Finally, humans still outperform even powerful VLMs overall. We identify three common error types in VLM reasoning about multimodal figurative phenomena: hallucination and incomplete or unsound reasoning.
7 Acknowledgments
We would like to thank the annotators for their work, as well as anonymous reviewers for productive discussion and feedback. This research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200005. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes notwithstanding any copyright annotation therein.
@@ 9 Limitations
Reference-based evaluation has fundamental flaws, such as not considering all possible explanations, which would be impossible to collect. However, current reference-free metrics for free-text rationales may still have flaws such as bias toward length or the evaluator LLM Stureborg et al. (2024); Raina et al. (2024); Huang et al. (2024); Chiang and Lee (2023); Wei et al. (2024). When evaluating textual explanations against these references, as is the case with any reference-based evaluation, there may also be a preference towards models which output text closer in distribution to the GPT-4 model. Because of that, it is important to utilize the data set in order to compare models other than the teacher model and pay more attention to the F1@0 scores, which represent simple classification scores and do not require the outputs to be similar in distribution. In terms of pure F1 score performance, GPT-4 underperforms the fine-tuned model, and performs very closely with Gemini and Claude that were not used to generate the data, with less than 2% difference (see column F1@0, Table 4). Although we showed a relatively high predictive power of automatic explanation scores to predict human judgments (see Section 5.3), future work may focus on increasing reliability of reference-based and reference-free textual explanation evaluation methods.
@@
Appendix A Details on Expert Verification
@@
Appendix B Dataset Statistics
@@
Appendix C API models Hyperparameters
C.1 Claude
@@ C.3 Gemini
Appendix D Fine-tuning Hyperparameters
@@
We utilize LoRA Hu et al. (2022) to fine-tune the models. We utilize the same hyperparameters for all fine-tunes outlined in Appendix D and use early stopping based on a V-FLUTE validation set to prevent overfitting. For evil and e-ViL+V-FLUTE we only fine-tuned for 2 epochs due to size of the e-ViL dataset and took the best checkpoint based on early stopping on V-FLUTE validation set. For eViL we only fine-tuned for 1 epoch to prevent overfitting. For VFLUTE, we trained for 3 epochs, and VFLUTE -Image for 10 epochs (to ensure performance does not increase even with larger number of epochs), for both we took the best checkpoint based on early stopping.
@@
Appendix E Prompts for LLMs
E.1 HAIVMET
E.1.1 One-shot Prompt for generating explanations
@@
E.2 IRFL
E.2.1 Zero-shot Prompt for generating explanations
@@
E.3 MuSE
E.3.1 Few-shot Prompt for generating opposite claims
@@
E.3.2 Zero-shot Prompt for Rephrasing
@@
E.4 MemeCap
E.4.1 Few-shot Prompt for generating entailing claims